OpenStudy (anonymous):

A construction company rented bins from a waste collection agency for a large-scale clean-up. The bins are in the shape of rectangular prisms. The dimensions of a bin are shown below. • Length – 9 feet • Width – 6 feet • Height – 4.5 feet How much waste material did the construction company collect if 8 such bins were filled to the brim with waste material? Answer 1,080 cubic feet 1,296 cubic feet 1,728 cubic feet 1,944 cubic feet

OpenStudy (anonymous):

Volume of bin = 9*6*4.5=243 cubic feet. Total waste material, 243*8= 1944 cubic feet.